# Tendon-Based Proprioception in an Anthropomorphic Underactuated Robotic Hand with Series Elastic Actuators

### v1

@@ III-C Tension Profile Characterization
When estimating joint angles from contact detection and relative stiffness from slope, it is critical to constrain the object, which is typically achieved through thumb opposition. If a finger contacts the object before opposition, object pushing may occur as in the nominal grasping case in Fig. 5(d), displacing the object and introducing errors in estimation. To prevent this, a strategy similar to that of [26] was adopted: the first contacting finger is paused, and actuation resumes only after all fingers have made contact. This contact-pause strategy ensures that the object retains its initial position.
@@
Contact detection from tension profile deviations is implemented using a sliding-window root-mean-square error (RMSE) method. The control loop operates at $1000\text{\,}\mathrm{Hz}\text{/}$, and RMSE is computed over a 50-sample window ($50\text{\,}\mathrm{ms}\text{/}$). Before computation, the window is offset-corrected by aligning its first sample to the corresponding reference value, ensuring consistent comparison with the reference profile. When the RMSE exceeds a predefined threshold $R_{\mathrm{prox}}$, contact is detected and assigned to the midpoint of the window. To identify the initial contact type, the global slope is evaluated $75\text{\,}\mathrm{ms}\text{/}$ after threshold crossing: values below $S_{\mathrm{load}}$ indicate proximal contact, while values above $S_{\mathrm{load}}$ indicate distal contact. If the initial contact is distal, the system proceeds directly to rigidity evaluation; if proximal, an additional RMSE-based search with threshold $R_{\mathrm{dist}}$ is conducted to detect distal contact before rigidity evaluation. The object is classified as rigid if any local slope, computed over a sliding window, exceeds $S_{\mathrm{rigid}}$; otherwise, it is deemed soft, and actuation is halted to regulate the grasping force, thereby preventing excessive deformation or fracture. To prevent misclassification or instability due to object stabilization, actuation is also terminated if the local slope falls below $S_{\mathrm{instab}}$ during loading. This rule-based strategy enables accurate estimation and stable grasp execution. Fig. 5(e) illustrates the entire grasping algorithm implemented for each finger.
III-D Finger Actuation Experiment
@@
To validate the accuracy of the modeling and the plausibility of the tension profile characterization, experiments were conducted using a PSF module. Fig. 6(a) shows the setup for the free-flexion cases. Two representative orientations were selected: one with the palm facing sideways, unaffected by gravity, and another with the palm facing downward, subject to gravity. The torsional spring preloads of the MCP and PIP joints were also adjusted in two combinations to examine different actuation sequences: a late-PIP condition (MCP $36\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$/PIP $32\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$) and an early-PIP condition (MCP $28\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$/PIP $12\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$). The results are shown in Fig. 6(b). Under the late-PIP condition, the joint angles and tension matched the modeling reference with high accuracy. Under the early-PIP condition, slight discrepancies were observed in the PIP joint, likely because the weaker preload caused two joints to actuate more simultaneously, making the motion susceptible to friction. In both conditions, gravitational effects were minimal, as the finger was lightweight with relatively high joint stiffness, making spring forces dominant. To prevent the ejection in underactuated fingers and to secure a larger workspace, the late-PIP condition was selected as the final parameter.
@@
Fig. 6(c) shows the setup for finger–object contact cases. Both adaptive and distal-only grasp types were considered with a fixture substituting for the role of thumb opposition. The test object was a $50\text{\,}\mathrm{mm}\text{/}$ diameter cylinder; the rigid one was 3D-printed in PA12 (3D HR PA12, HP Inc., USA), and the soft ones were printed in TPU (VarioShore TPU, colorFabb, Netherlands) with varying infill densities. A gyroid infill pattern, which is nearly isotropic [27], was selected. For this experiment, actuation was not halted as in Fig. 5(e) but instead uniformly increased to $40\text{\,}\mathrm{N}\text{/}$ to observe slope variations by deformation during loading. The RMSE thresholds for contact detection were set to $R_{\mathrm{prox}}=$0.4\text{\,}\mathrm{N}\text{/}$\$ and $R_{\mathrm{dist}}=$1.0\text{\,}\mathrm{N}\text{/}$\$, with the loading slope threshold set to $S_{\mathrm{load}}=$2\text{\,}\mathrm{N}\text{/}\mathrm{mm}$\$.
@@
Fig. 6(d) presents the modeled reference and experimental joint angles and tension profiles with RMSE and slopes. The reference profiles for joint angles and tension were generated for the case of a rigid object. In both grasp types, the profiles clearly exhibited the distinct proximal and distal contact patterns characterized in Fig. 5(a), and the slopes varied noticeably with stiffness level as in Fig. 5(b). In both grasp types, the slope for the rigid object was slightly lower than the reference, likely due to deformation of the silicone fingertip. Since this deformation consistently occurs at every contact, the experimentally measured rigid slope can serve as a practical reference. Between the two slope metrics, the local and global slopes exhibited similar tendencies, but the global slope proved more stable and reliable for distinguishing relative stiffness levels. In the adaptive grasp, rigid objects caused the MCP angle to remain fixed after proximal contact, whereas soft objects allowed the MCP angle to continue increasing slightly along with the PIP joint. This difference makes estimation more challenging for soft objects. Nevertheless, the rule-based contact detection scheme naturally accounts for such interaction. Since the weak resistance of soft objects produces only a small deviation from the reference profile, it delays the threshold crossing. These results yielded larger estimated MCP angles and smaller PIP angles, thereby mitigating potential errors and preserving consistency with the actual grasp dynamics. Meanwhile, in the distal-only grasp—where the interaction is simpler than in the adaptive case—the contact timing was consistent regardless of object stiffness.
IV Grasp Functionality Implementation
IV-A Disturbance Detection with Compliance
IV-B Grasp Posture Estimation
@@
Figure 10: Blind grasping and object recognition demonstration. (a) Sample Objects with varying size, shape, and stiffness. (b) Task setup with exploration trajectory. (c) Task completion with object recognition result. (d) Execution of approach and grasp after palmar and dorsal contact.
The joint angle estimation of individual fingers was extended to reconstruct the overall grasp posture. Three irregularly shaped YCB objects [28] were selected: a power drill, a wine glass, and a plastic pear. The drill and glass correspond to adaptive grasps, while the pear corresponds to a distal-only grasp. Actual joint angles were measured with a motion capture system (OptiTrack, NaturalPoint Inc., USA). Two reflective markers were attached to each phalanx, and joint angles were computed from changes in marker vectors between initial and final configurations. As shown in Fig. 8, the estimated posture from tension profiles was visualized on a CAD model and compared with the actual posture. The average error per joint angle was $3.0\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$ for the drill and glass and $2.1\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$ for the pear, indicating high overall accuracy.
IV-C Safe Grasp Force Regulation for Deformable Objects
@@
IV-D Blind Grasping with Object Recognition
@@
A blind grasping task with object recognition was conducted to demonstrate the integrated proprioceptive functionality. This task was inspired by the way humans sweep their hands across a table to locate and grasp objects without vision. As shown in Fig. 10(b), the hand mounted on a robot manipulator (RB5-850, Rainbow Robotics, Korea) explored a defined search area by executing three consecutive sweeps—palmar, dorsal, and palmar. Before exploration, the fingers were flexed slightly for $50\text{\,}\mathrm{ms}\text{/}$ to preload tension, thereby enabling bidirectional disturbance detection. Once contact was detected, the hand executed a predefined approach and completed the grasp, with dorsal contacts leading to an approach from behind the object as shown in Fig. 10(d).
@@
Seventeen objects with varying size, shape, and stiffness were used, including cylinders, cones, and spheres in rigid and soft forms, as shown in Fig. 10(a). All objects were fabricated with the same material as in Fig. 6(c). To prevent rolling or slipping during exploration, spheres were mounted on support bases, and soft objects had sandpaper attached to their bottoms. For classification, a simple k-nearest neighbors (k-NN) model ($k=3$, Euclidean distance) was employed as a proof-of-concept demonstration. Object geometry (size and shape) was recognized using a kinematics-based approach [9, 29], while stiffness recognition was subsequently performed within each geometry category by comparing the average global slope across five fingers. The feature vector of the training dataset consisted of nine joint angles and the averaged global slope, collected from ten grasp trials per object. To enhance robustness against variability in the approach motion, the object positions were slightly varied during training. As depicted in Fig. 10(c), the system successfully identified unknown objects and displayed the recognition result, with the complete process provided in the supplementary video.
V Discussion & Conclusion
This study presented an integrated system for tendon-based proprioception in an anthropomorphic underactuated hand, addressing hardware design for compact actuation as well as the analysis and application of proprioceptive tension profiles. To this aim, the MPC-SEA was developed as a compact module providing reliable tension sensing and enabling seamless embedding of proprioception within the hand. With an energy-based model, the tension profile serves as an informative grasp-state indicator, enabling comprehensive hand–object interactions throughout the entire grasp process. During reaching, contact detection enables joint angle estimation; during loading, tendon excursion associated with deformation supports relative stiffness evaluation; during holding, finger configuration changes reflected in tension indicate external disturbances. Additionally, these interaction variables offer consistent and quantitative grasp-state information, supporting applicability across different tasks and hardware settings. Experiments validated the sensing performance of the MPC-SEA and confirmed the characterization of tension profiles for hand–object interaction. In addition, demonstrations highlighted key functionalities including grasp posture estimation, safe handling of deformable objects, and proprioceptive recognition in blind grasping.

### v2

@@ III-B Tension Profile Characterization
III-C Grasping Strategy and Algorithm for State Estimation
Figure 7: Grasping strategy and algorithm. (a) Contact-pause strategy to prevent object displacement and maintain the initial location. (b) Rule-based grasping algorithm.
@@ III-C Grasping Strategy and Algorithm for State Estimation
When estimating joint angles from contact detection and relative stiffness from slope, it is critical to constrain the object, which is typically achieved through thumb opposition. If a finger contacts the object before opposition, object pushing may occur as in the nominal grasping case in Fig. 7(a), displacing the object and introducing errors in estimation. To prevent this, a strategy similar to that of [23] was adopted: the first contacting finger is paused, and actuation resumes only after all fingers have made contact. This contact-pause strategy ensures that the object retains its initial position.
@@
Contact detection from tension profile deviations is implemented using a sliding-window root-mean-square error (RMSE) method. The control loop operates at $1\text{\,}\mathrm{kHz}\text{/}$, and RMSE is computed over a 50-sample window ($50\text{\,}\mathrm{ms}\text{/}$). Before computation, the window is offset-corrected by aligning its first sample to the corresponding reference value, ensuring consistent comparison with the reference profile. When the RMSE exceeds a predefined threshold $R_{\mathrm{prox}}$, contact is detected and assigned to the midpoint of the window. To identify the initial contact type, the global slope is evaluated $75\text{\,}\mathrm{ms}\text{/}$ after threshold crossing: values below $S_{\mathrm{load}}$ indicate proximal contact, while values above $S_{\mathrm{load}}$ indicate distal contact. If the initial contact is distal, the system proceeds directly to rigidity evaluation; if proximal, an additional RMSE-based search with threshold $R_{\mathrm{dist}}$ is conducted to detect distal contact before rigidity evaluation. The object is classified as rigid if any local slope, computed over a sliding window, exceeds $S_{\mathrm{rigid}}$; otherwise, it is deemed soft, and actuation is halted to regulate the grasping force, thereby preventing excessive deformation or fracture. To prevent misclassification or instability due to object stabilization, actuation is also terminated if the local slope falls below $S_{\mathrm{instab}}$ during loading. This rule-based strategy enables accurate estimation and stable grasp execution. Fig. 7(b) illustrates the entire grasping algorithm implemented for each finger.
IV Experimental Validations
IV-A MPC-SEA Experiment
To evaluate the accuracy and reliability of the MPC-SEA’s tension sensing, a cyclic loading test was conducted. As shown in Fig. 8(a), the cable of a fixed MPC-SEA was connected to a load cell ( [^3] FDX, KTOYO, Korea). Each cycle consisted of loading from $0\text{\,}\mathrm{N}\text{/}$ to $45\text{\,}\mathrm{N}\text{/}$ ($\approx$90% of maximum tension) and then unloading back to $0\text{\,}\mathrm{N}\text{/}$. The tension was recorded over loading, which is relevant to grasping, and compared with the reference tension measured by the load cell. Before the test, a MPC-SEA was calibrated to compensate for slight parameter variations among components, such as springs and potentiometers, as well as minor offsets introduced by the potentiometer’s placement during assembly. The calibration was performed using data from the 100 cycles, which were fit with a first-order linear model via the least squares method to determine the slope and intercept of (3). After the calibration process, a total of 50,000 cycles was performed.
IV-B Finger Actuation Experiment
IV-B1 Free-Flexion Case
To validate the accuracy of the modeling, experiments were conducted using a PSF module. Fig. 9(a) shows the setup for the free-flexion cases. Two representative orientations were selected: one with the palm facing sideways, unaffected by gravity, and another with the palm facing downward, subject to gravity. The torsional spring preloads of the MCP and PIP joints were also adjusted in two combinations to examine different actuation sequences: a late-PIP condition (MCP $36\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$/PIP $32\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$) and an early-PIP condition (MCP $28\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$/PIP $12\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$). The actuation was performed by applying the maximum PWM command, resulting in a closing time of $1.3\text{\,}\mathrm{s}\text{/}$ at the highest motor speed. The real-time tension was sensed through the SEA, while the actual joint angles were measured with a motion capture system (OptiTrack, NaturalPoint Inc., USA). Two reflective markers were attached to each phalanx, and joint angles were computed from changes in marker vectors between initial and final configurations.
@@ IV-B1 Free-Flexion Case
Each experiment was repeated five times, and the average results are shown in Fig. 9(b). Under the late-PIP condition, the joint angles and tension matched the modeling reference with high accuracy. Under the early-PIP condition, slight discrepancies were observed in the PIP joint, likely because the weaker preload caused the two joints to actuate more simultaneously, making the motion more susceptible to friction. In both conditions, gravitational effects were minimal, as the finger was lightweight with relatively high joint stiffness, making spring forces dominant. To prevent ejection in underactuated fingers and to secure a larger workspace, the late-PIP condition was selected as the final parameter. Although the model is based on a quasi-static assumption, the experimental results demonstrate that it remains valid even under dynamic actuation. Tendon-based proprioception allows for a simple finger structure without additional sensors, yielding a lightweight finger of only about $28.6\text{\,}\mathrm{g}\text{/}$; thus, inertial effects are negligible, and the quasi-static assumption remains well justified.
IV-B2 Object Contact Case
@@
To validate the plausibility of the tension-profile characterization, the experimental setup for finger–object contact cases was configured as shown in Fig. 9(c). Both adaptive and distal-only grasp types were considered, with a fixture substituting for the role of thumb opposition. The test object was a $50\text{\,}\mathrm{mm}\text{/}$ diameter cylinder; the rigid one was 3D-printed in PA12 (3D HR PA12, HP Inc., USA), and the soft ones were printed in TPU (VarioShore TPU, colorFabb, Netherlands) with varying infill densities. A gyroid infill pattern, which is nearly isotropic [24], was selected. For this experiment, actuation was not halted as in Fig. 7(b) but instead uniformly increased to $40\text{\,}\mathrm{N}\text{/}$ to observe slope variations by deformation during loading. The RMSE thresholds for contact detection were set to $R_{\mathrm{prox}}=$0.4\text{\,}\mathrm{N}\text{/}$\$ and $R_{\mathrm{dist}}=$1.0\text{\,}\mathrm{N}\text{/}$\$, with the loading slope threshold set to $S_{\mathrm{load}}=$2\text{\,}\mathrm{N}\text{/}\mathrm{mm}$\$.
@@
Fig. 9(d) presents the modeled reference and experimental joint angles and tension profiles with RMSE and slopes, and Table III summarizes the estimated joint angles with the actual joint angles measurements and the relative stiffness values for different grasp conditions. The reference profiles for joint angles and tension were generated for the case of a rigid object, obtained by solving the optimization problem in (12) while constraining the $\mathrm{RoM}_{j}$ to the angles expected at object contact. The profiles clearly exhibited the distinct proximal and distal contact patterns characterized in Fig. 6(a), and the slopes varied noticeably with stiffness level as in Fig. 6(b). In both grasp types, the slope for the rigid object was slightly lower than the reference, likely due to deformation of the silicone fingertip. Since this deformation consistently occurs at every contact, the experimentally measured rigid slope can serve as a practical baseline. Between the two slope metrics, the local and global slopes exhibited similar tendencies, but the global slope proved more stable and reliable for distinguishing relative stiffness levels.
TABLE III: Joint Angle and Relative Stiffness Estimation under Different Grasp Conditions
Parameter Symbol Adaptive Grasp Distal-Only Grasp
Rigid Soft 12% Soft 6% Soft 3% Rigid Soft 12% Soft 6% Soft 3%
MCP Joint Angle $\theta_{1,\mathrm{est}}$ [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 46.10 47.28 49.13 56.77 58.20 58.76 58.84 58.68
$\theta_{1,\mathrm{act}}$ [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 44.20 46.49 49.30 53.82 57.64 58.31 58.48 57.10
error1 [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 1.90 0.79 -0.17 2.95 0.56 0.45 0.36 1.58
PIP Joint Angle $\theta_{2,\mathrm{est}}$ [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 56.42 55.12 51.47 43.44 15.07 15.07 15.07 15.07
$\theta_{2,\mathrm{act}}$ [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 54.80 50.92 43.91 37.54 16.17 15.88 16.10 15.90
error2 [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 1.62 4.20 7.56 5.90 -1.10 -0.81 -1.03 -0.83
Relative Object Stiffness $k_{\mathrm{rel}}$ [N/$m$] 4.26 3.63 2.96 2.65 4.27 3.95 3.75 2.45
In the adaptive grasp, rigid objects caused the MCP angle to remain fixed after proximal contact, whereas soft objects allowed the MCP angle to continue increasing slightly along with the PIP joint. Namely, tendon excursion that would have contributed solely to PIP flexion after proximal contact in the rigid object case also contributes to MCP flexion when interacting with soft objects, making joint angle estimation more challenging for soft objects. Nevertheless, the rule-based contact detection scheme naturally accounts for such interaction. Since the weak resistance of soft objects produces a smaller deviation from the reference profile, it delays the proximal threshold ($R_{\mathrm{prox}}$) crossing. These results yielded larger estimated MCP angles and smaller PIP angles, thereby mitigating potential errors and preserving consistency with the actual grasp dynamics. As shown in Table III, the MCP joint-angle estimation errors were all within $3\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$. In contrast, the PIP estimation errors were relatively larger for softer objects, mainly because the deviation at distal contact is smaller than in the rigid object case, which delays the distal threshold ($R_{\mathrm{dist}}$) crossing and leads to slightly overestimated PIP angles. While the estimation scheme is primarily tailored to rigid object behavior, it nonetheless maintains acceptable performance when applied to soft objects as well.
Meanwhile, in the distal-only grasp, where the interaction is simpler than in the adaptive grasp, the contact excursion remained nearly unchanged across different object stiffnesses. Therefore, the estimated joint angles remained similar regardless of object stiffness, and the estimation errors were consistently very small, indicating that accurate joint angle estimation was achieved.
Figure 10: Heatmap analysis of joint angle estimation errors across threshold $R_{\mathrm{prox}}$ and $R_{\mathrm{dist}}$.
In addition, to examine the influence of the threshold value selection in the rule-based algorithm, the joint angle estimation errors were evaluated while varying $R_{\mathrm{prox}}$ and $R_{\mathrm{dist}}$ within the ranges 0.1–0.7 $\text{\,}\mathrm{N}\text{/}$ and 0.7–1.3 $\text{\,}\mathrm{N}\text{/}$, respectively, as shown in Fig. 10. During this analysis, $S_{\mathrm{load}}$ was fixed at $2\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}\mathrm{m}$, since this threshold was found to be insensitive; varying it between 1–3 did not affect the estimation results. Fig. 10 shows that the PIP angle is generally more sensitive to threshold selection than the MCP angle. When $R_{\mathrm{prox}}$ is too small, the estimator becomes sensitive to noise and may falsely detect contact during free flexion. In contrast, if $R_{\mathrm{prox}}$ is set too high, the proximal contact deviation for soft objects may not reach the threshold, increasing estimation error. $R_{\mathrm{dist}}$ has only a minor influence, as the distal contact produces a clear inflection in the tension profile. Sensitivity also increases with softer objects, whose tension profiles deviate more from the modeled reference. This effect is most pronounced for the soft object of 3% infill, where improper threshold selection can lead to noticeable errors. These observations collectively indicate that while the proposed estimation scheme is broadly robust, careful selection of $R_{\mathrm{prox}}$ is important, particularly for highly soft objects.
V Grasp Functionality Demonstration
V-A Grasp Posture Reconstruction
@@
The joint angle estimation of individual fingers was extended to reconstruct the overall grasp posture. Three irregularly shaped YCB objects [25] were selected: a power drill, a wine glass, and a plastic pear. The drill and glass correspond to adaptive grasps, while the pear corresponds to a distal-only grasp. Actual joint angles were measured using the same motion capture procedure described in Section 9. As shown in Fig. 11, the estimated posture from tension profiles was visualized on a CAD model and compared with the actual posture. The average error per joint angle was $3.0\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$ for the drill and glass and $2.1\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$ for the pear, indicating high overall accuracy.
V-B Safe Grasp Force Regulation for Deformable Objects
@@
V-C Disturbance Detection with Compliance
V-D Blind Grasping with Object Recognition
Figure 14: Blind grasping with object recognition demonstration. (a) Sample Objects with varying size, shape, and stiffness. (b) Task setup with exploration trajectory. (c) Task completion with object recognition result. (d) Execution of approach and grasp after palmar and dorsal contact.
@@ V-D Blind Grasping with Object Recognition
A blind grasping task with object recognition was conducted to demonstrate the integrated proprioceptive functionality. This task was inspired by the way humans sweep their hands across a table to locate and grasp objects without vision. The demonstration integrates three key capabilities: object localization through disturbance detection, autonomous grasp force regulation for safely handling soft objects, and object recognition based on grasp posture reconstruction and stiffness classification. As shown in Fig. 14(b), the hand mounted on a robot manipulator (RB5-850, Rainbow Robotics, Korea) explored a defined search area by executing three consecutive sweeps—palmar, dorsal, and palmar. Before exploration, the fingers were flexed slightly for $50\text{\,}\mathrm{ms}\text{/}$ to preload tension, thereby enabling bidirectional disturbance detection. Once contact was detected, the hand executed a predefined approach and completed the grasp, with dorsal contacts leading to an approach from behind the object as shown in Fig. 14(d).
@@
Seventeen objects with varying size, shape, and stiffness were used, including cylinders, cones, and spheres in rigid and soft forms, as shown in Fig. 14(a). All objects were fabricated with the same material as in Fig. 9(c). To prevent rolling or slipping during exploration, spheres were mounted on support bases, and soft objects had sandpaper attached to their bottoms. For object recognition, a simple k-nearest neighbors (k-NN) model ($k=3$, Euclidean distance) was employed as a proof-of-concept demonstration. Object geometry (size and shape) was recognized using a kinematics-based approach that utilizes grasp posture [8, 26], while stiffness classification was subsequently performed within each geometry category by comparing the average relative stiffness across five fingers. The feature vector of the training dataset consisted of nine joint angles and the averaged relative stiffness, collected from ten grasp trials per object. To enhance robustness against variability in the approach motion, the object positions were slightly varied during training. As depicted in Fig. 14(c), the system successfully identified unknown objects and displayed the recognition result, with the complete process provided in the supplementary video.
The confusion matrices in Fig. 15 show the recognition performance with 5-fold cross-validation. Geometry classification achieved an accuracy of 95.3%. Stiffness classification for cones and spheres also showed high accuracy, whereas cylinders exhibited slightly lower performance, achieving around 80% accuracy. Nevertheless, the overall demonstration results confirm that the proposed estimation scheme remains effective in identifying both object geometry and stiffness using only tendon-based proprioception, without reliance on vision or tactile sensors.
Figure 15: Confusion matrices for object recognition k-NN model. (a) Geometry. (b)–(h) Stiffness in each geometry category.
VI Discussion & Conclusion
This study presented an integrated system for tendon-based proprioception in an anthropomorphic underactuated hand. The MPC-SEA was developed as a compact sensing–actuation module and integrated with the sensorless fingers and on-board electronics to form the complete system. By utilizing an energy-based finger model, the tension profile serves as an informative grasp state indicator, enabling comprehensive interpretation of hand–object interaction throughout the entire grasp sequence. The proposed grasp state estimation framework is also scalable to general tendon-driven underactuated finger designs with quantitative grasp state information, as long as the joint stiffness is well defined and the system energy can be modeled. To summarize, the main contributions are as follows: 1) A generalizable grasp state estimation framework for underactuated hands that infers multiple grasp states from only a single proprioceptive tension signal, using energy-based finger modeling and an efficient rule-based algorithm. 2) Development of the MPC-SEA module distinguished by its high sensing accuracy and reliability with full integrability into compact robotic hand structures. 3) Experimental validation demonstrating the efficacy of the proposed framework and extensive demonstrations showing practical grasp functionalities.